# EDGE: Enhanced Grounded GUI Understanding with Enriched Multi-Granularity Synthetic Data

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: THE WEB CONFERENCE 2025; April 28, 2025; Sydney, New South Wales, Australiaisbn: 978-1-4503-XXXX-X/18/06submissionid: 2903copyright: none
@@ 5.4. Ablation Study
We further present two toy examples in Figure 6 to display the different outputs of the full EDGE and its ablated variant. The full EDGE perfectly answered the questions in the first screenshot, going beyond repeating the heading. In the second one, it successfully distinguished between the product itself and the button for adding it to the cart. These examples provide intuitive demonstrations of how training on advanced tasks can enhance the model. Additional demonstration examples are provided in Appendix C.
@@ 7. Conclusion
Appendix A Annotation Script
@@
Next, we determine the type of each visible element based on attributes like tag, role, and computed styles, covering common elements found on webpages, such as text, code, images, icons, buttons, hyperlinks, and inputs. Then we designed a set of integration rules based on the type, content, and children of an element to decide whether it represents the smallest, complete semantic unit visually. For example (illustrated in Figure 7), buttons often consist of borders, text, and icons visually. Traditional methods tend to recognize the text and icon separately, leading to incomplete semantic understanding. In contrast, we treat the entire button as a single element and use the outermost border to determine its position. This approach aligns more closely with human visual perception, providing a more cohesive semantic representation.
@@ B.1. Training Samples
When rendering the webpage, we select from 16 various screen sizes to simulate the presentation of webpages on desktop, tablet and mobile devices. The input images of the training samples are raw screenshots. The dotted box annotations in the previous images are only for demonstration purposes.
We use two question forms in grounding tasks: predicting the bounding box $(x1,y1,x2,y2)$ or its center point $(x,y)$, with the specific choice explicitly indicated in the input. As shown in Figure8, for element-level tasks, we use a multi-turn conversation format, which is not only more sample efficient, but also bring better experimental results. Page-level tasks still occupy a single sample.
@@
Appendix C More Inference Results
More reasoning results are shown in Figure 9. Our model supports understanding and answering user operation instructions in a natural, human-like manner.
@@ Appendix C More Inference Results
Figure 9. Examples of inference results of EDGE on web, desktop, and mobile environments are shown. The output bounding box is marked with a red rectangular marks.

### v2

@@ Abstract.
conference: Under Review; 2024; copyright: none
@@ 5.4. Ablation Study
We further present two toy examples in Figure 6 to display the different outputs of the full EDGE and its ablated variant. The full EDGE perfectly answered the questions in the first screenshot, going beyond repeating the heading. In the second one, it successfully distinguished between the product itself and the button for adding it to the cart. These examples provide intuitive demonstrations of how training on advanced tasks can enhance the model. Additional demonstration examples are provided in Appendix D.
@@ 7. Conclusion
Appendix A The Annotation Script
@@
Next, we determine the type of each visible element based on attributes like tag, role, and computed styles, covering common elements found on webpages, such as text, code, images, icons, buttons, hyperlinks, and inputs. Then we designed a set of integration rules based on the type, content, and children of an element to decide whether it represents the smallest, complete semantic unit visually. For example (illustrated in Figure 7), buttons often consist of borders, text, and icons visually. Traditional methods tend to recognize the text and icon separately, leading to incomplete semantic understanding. In contrast, we treat the entire button as a single element and use the outermost border to determine its position. This approach aligns more closely with human visual perception, providing a more cohesive semantic representation. Figure 8 shows the more webpage annotations results.
Figure 8. More webpage annotations results.
[…]
More webpage annotations results.
In practice, to avoid samples being concentrated in the header of webpages, we dynamically determine the number and location of annotations based on the length of webpages. Specifically, if the webpage is short and is completely contained in the current viewport, it will be only annotated once; if the webpage is longer, it will be first annotated at the header, and then we scroll to its bottom to annotated it again; if the webpage is very long, it will be annotated once more in the middle.
@@ B.1. Training Samples
When rendering webpages, we choose from 16 pre-selected common screen sizes to simulate the presentation of webpages on desktop, tablet and mobile devices.
For elementary tasks, the input images of the training samples are raw screenshots, and the dotted box annotations in the previous images are only for demonstration purposes. We use two question forms in grounding tasks: predicting the bounding box $(x1,y1,x2,y2)$ or its center point $(x,y)$, with the specific choice explicitly indicated in the input. As shown in Figure9, for element-level tasks, we use a multi-turn conversation format, where a sample consists of which is not only more sample efficient, but also bring better experimental results. Page-level tasks still occupy a single sample.
Since the encoding length of the Monkey (Li et al., 2024b) for an image exceeds half of the total sequence length, each training sample contains at most one image. We adopt different sample formats for element-level tasks and page-level tasks. As shown in the Figure  9, for element-level tasks like OCR and grounding, we use a multi-turn conversation format, where each training sample contains multiple independent question-answering instances of a single screenshot on the same task. This setting brings a higher density of supervision signals, makes full use of the calculation of each forward propagation, and can improve the efficiency of model training. Page-level tasks like captioning still occupy a single sample. Page-level question-answering tasks such as captioning still take a single sample. Advanced tasks share a similar division, with element-level the conversation intention task using a multi-turn conversation setting, while the page-level detailed description and function inference use a single question-answer format.
@@
[…]
Examples of training samples on the homepage of Google.
B.2. Question Templates for Elementary Tasks
For single-round QA tasks, we first manually write three question templates for each task, and then prompted ChatGPT to generate more templates that are richer and more various in natural language features such as vocabulary, syntax, and sentiment, to simulate the query input by various users with different prompting habits. For multi-turn QA tasks such as OCR and grounding, as well as all of mobile screen tasks on the Rico dataset, we directly use the multi-turn question templates provided by SeeClick (Cheng et al., 2024), which typically have dozens of templates for each task. Note that as shown in Figure 9 in the element-level multi-turn conversation, the ”template” we refer to here is the task description. For example, for the grounding task, we first input the task description, followed by multiple elementary questions and answers, i.e., the description and the location of the on-screen elements. As for the page-level tasks, templates are the sepcific questions. Figure 10 shows some templates for grounding, OCR, captioning, and icon describing tasks.
B.3. Prompts for Advanced Tasks
We use Claude-3.5-Sonnet-20240620 to generate advanced task QA data. According to preliminary manual judgment, the quality of QA pairs generated by Claude-3.5 is slightly higher than those by GPT-4V. Following Ferret-UI (You et al., 2024), the textual prompts we input into Claude include shared prompts, task prompts, screen annotations, and QA examples (only for conversation intention). On this basis, webpage screenshots annotated with the Set-of-Mark (Yang et al., 2023b) visual prompting method are used as visual input to help the model capture information and semantics that the automatic annotation may miss. Human evaluation shows that the Set-of-Mark prompted screenshots input effectively utilizes the powerful visual understanding ability of proprietary models, which can not only correct the semantics that are omitted or mislabeled by the annotation script, but also reduce the model’s hallucinations. Because we require the model to output the annotation number corresponding to each element, which forces it to select the annotated elements, thereby using accurate positions instead of the inaccurate self-generated ones.
An example of the conversation intention task is displayed in Figure 9. Examples of the shared prompt and the task prompt for the three advanced tasks are shown in Figure 11.
Figure 10. Question templates for some elementary tasks.
[…]
Question templates for some elementary tasks.
Figure 11. Prompts for advanced tasks data synthesizing using Claude-3.5. Some prompts are slightly simplified, retaining the core logic and omitting the cumbersome details.
[…]
Prompts for advanced tasks data synthesizing using Claude-3.5. Some prompts are slightly simplified, retaining the core logic and omitting the cumbersome details.
Appendix C Evaluation Details
C.1. VisualWebBench
There are seven tasks in VisualWebBench (Liu et al., 2024d): Caption, WebQA, HeadOCR, Element OCR, Element Ground, Action Prediction, and Action Ground.
The Caption task requires models to predict the content attribute of the tag <meta name="description" content="xxx"> in the HTML header. The WebQA task involves questions and answers about various contents on webpages. HeadOCR requires the model to output the content of the most eye-catching heading on the webpage. The Element OCR task gives the bounding box of an element and marks the element with a red rectangular box in the screenshot, requiring the model to extract the text content of the element. Element Ground provides the text of an element, marks multiple elements with numbered red rectangular box in the screenshot, and requires the model to select the correct number, that is, the number corresponding to the element containing the given text. In the Action Prediction task, the model choose the most likely description of the webpage that will be jumped to after clicking on an element. In the Action Ground task, the model selects the element to be clicked based on the given user instructions and the screenshot with multiple elements marked with red boxes, also in the form of multiple choices.
When evaluating the GUI understanding and interaction capabilities of LVLMs on VisualWebBench (Liu et al., 2024d), we adopt the prompts provided by VisualWebBench by default. The exceptions are that for the evaluation of EDGE on the Element Ground and Action Ground tasks, we use our own prompts, requiring the model to directly predict the coordinate point instead of the default multiple choice setting.
We also slightly modified the prompt of the Caption task to describe the target of the task in a more humane way. In contrast, the original prompt directly requires the model to output <meta name="description" content="YOUR ANSWER">, which is too formalized. Intuitively, neither humans nor GUI interaction models need to understand the HTML structure and content behind the webpage.
For the above three tasks, our modified prompts are shown in Figure 12.
Figure 12. Modified task prompts in VisualWebBench.
[…]
Modified task prompts in VisualWebBench.
C.2. ScreenSpot
Each sample in ScreenSpot (Cheng et al., 2024) is an action grounding task. Experimental results show that the point prediction of each model is slightly better than the bounding box prediction. Therefore, we use the same prompt as the Action Ground task of VisualWebBench, as shown in Figure 12.
C.3. Downstream Agent Benchmarks
We use the same dataset and evaluation settings as in SeeClick (Cheng et al., 2024) for fine-tuning and evaluating EDGE on the three downstream agent benchmarks, namely MiniWob (Shi et al., 2017), AITW (Rawles et al., 2024), and Mind2Web (Deng et al., 2024). Specifically, EDGE does not use any explicit agent framework, but simply concatenates instructions and the previous actions as text prompts, and then directly generates the next action. SeeClick has defined a set of mappings from textual actions to numerical action ids, and the fine-tuning process is to make the model familiar with the mapping between the action ids of the corresponding benchmark and the real action in natural language. Figure 13 shows the prompt format used by EDGE in the fine-tuning and evaluation of these three agent benchmarks, which is exactly the same as that shown in the appendix of the SeeClick paper (Cheng et al., 2024).
Figure 13. The prompt format for fine-tuning and evaluation in the three downstream agent benchmarks.
[…]
The prompt format for fine-tuning and evaluation in the three downstream agent benchmarks.
Appendix D More Inference Results
More reasoning results are shown in Figure 14. Here we ask the model to output the predicted bounding box of the target element instead of a single point to improve interpretability. The output bounding boxes are marked with red rectangular boxes in the screenshots.
The template-based elementary tasks enables the model to accurately locate elements, while after training on free-form high-level tasks, our model supports understanding and answering user operation instructions in a natural, human-like manner. With the help of LVLM’s native natural language understanding and generation capabilities, EDGE implicitly completes the process of intent recognition, action grounding, and element grounding, just like a human assistant talking to you.
@@ Appendix D More Inference Results
Figure 14. Examples of inference results of EDGE on web, desktop, and mobile environments are shown. The output bounding box is marked with a red rectangle.
[…]
Examples of inference results of EDGE on web, desktop, and mobile environments are shown. The output bounding box is marked with a red rectangle.